Solve: 3m^2+4=4m Thank You!
Answer by stanbon(75887)

x = 3y
x + 10y = 13
-----
Rearrange:
x - 3y = 0
x + 10y = 13
----
Subtract to eliminate the "x" terms then solve for "y":
13y = 13
y = 1
-----
Substitute into x = 3y to solve for "x":
x = 3y
x = 3*1
x = 3
---------
Cheers,
Stan H.